OpenStudy (anonymous):

s=80÷(1−(−1/2))= ? help

OpenStudy (lalaly):

80÷3/2 = 80* 2/3 = 160/3

OpenStudy (anonymous):

oh ok thanks'

OpenStudy (lalaly):

:)

OpenStudy (anonymous):

s=-80÷(1−(−1/2))= what about this would it be the same except negative

OpenStudy (anonymous):

S=160

OpenStudy (lalaly):

yeh the samee but with - sign

OpenStudy (anonymous):

ok thanks

OpenStudy (lalaly):

ur welcome

OpenStudy (anonymous):

WHO'S RIGHT?

OpenStudy (lalaly):

this tells u whos right